# LITMUS: Bayesian Lag Recovery in Reverberation Mapping with Fast Differentiable Models

### v1

@@ 3 Existing Methods
Reverberation mapping has progressed markedly over its decades long history, both in our understanding of the physics and of the best-practice in constraining the lags. Lag recovery methods can be broadly grouped into two categories: the Bayesian parametric methods which make use of our understanding of the AGN signal statistics, and the frequentist non-parametric methods which aim to be as agnostic as possible about the signal, using only flexible means of interpolation and measures of goodness of fit. In this section we focus primarily on the parametric family (Section 3.1), as this is the basis of the main fitting methods of LITMUS, but offer an overview of some popular methods / packages of non-parametric fitting as well (Section 3.2).
3.1 Parametric Methods
The main focus of this paper is the class of parametric methods that leverage our understanding of AGN signal statistics to work within the framework of a full Bayesian generative model. The core principle of these methods is to take the description of the AGN variability as a Gaussian process (discussed in detail in Section 2), for which there is a closed form and (somewhat) easily evaluable likelihood function.
@@ 3.1 Parametric Methods
3.1.1 JAVELIN
JAVELIN is the longest standing of the parametric fitting methods and has the distinction of being the most widely used. JAVELIN (JAVELIN-Zu_2010) is a successor to the FORTRAN-based SPEAR (Zu_2011_JAVELINpaper). It models the AGN as a damped random walk (following the covariance functions outlined in Section 2) and models the response smoothing with a “top-hat” smoothing function:
@@ 3.1.1 JAVELIN
Noting that the exact form of smoothing tends to have little impact on the lag recovery (OZDES-Yu_2019) it is common-place when using JAVELIN to fix the smoothing scale, $b$, to some fixed value shorter than the observation cadence (e.g. OzDES-Penton_2021), a choice found to have little impact on lag recovery in OzDES-like observational surveys (OZDES-Yu_2019).
@@
3.1.2 CREAM
CREAM (Starkey_2015_CREAM) is similar to JAVELIN in its modelling of the AGN light curve as a stochastic process and using Bayesian fitting, but differs in two key respects. Firstly, instead of fitting for the parameters of the observed signals it instead fits for physical properties of the AGN (e.g. mass, accretion rate etc.). Secondly it models the light curves by fitting the phase and amplitude of a series of Fourier components and, unlike JAVELIN, does not include the covariance of the observations in its likelihood function, instead using a $\chi^{2}$ goodness of fit loss function.
A novel feature of CREAM is that it can be set to rescale the error bars of measurements from different telescopes as a part of the MCMC fitting in an internal calibration process (starkey_2017_creamerrorscaling). CREAM is used as the primary lag recovery method by some SDSS releases (SDSS-Grier_2017; CIV_Grier_2019), and is included in the reported results of their final data release (SDSS-Shen_2023).
3.2 Non-Parametric Methods
Parametric methods offer the most complete way to perform lag recovery, they are also more complicated and computationally arduous. Aside from costing sheer time to compute, this also opens room for numerical errors (see discussion of the aliasing mixing problem in Section 2.1). For this reason, the less rigorous but more exhaustive family of non-parametric methods, which make as few assumptions about the underlying signal properties as possible, are still actively used in conjunction with or in preference over their more statistically complete parametric cousins.
Such non-parametric methods are valuable in that their lack of loose commitment towards a particular signal model lends them a flexibility and vagueness that absorbs our uncertainty about the AGN signal properties. Such models tend to also be numerically inexpensive, allowing exhaustive searches of their respective parameter spaces forcing their way past potential numerical challenges. Though LITMUS is primarily a parametric method, we include these here for the sake of completeness and comparison.
3.2.1 PyCCF
@@
The ICCF method has been found to agree with more rigorous models like JAVELIN to within statistical bounds, though with higher reported uncertainties (OZDES-Yu_2019). Owing to its numerical robustness and low computational overhead, it is still used as a validation tool by both OzDES (OzDES-Malik_2023; OzDES-Yu_2023; OzDES-Penton_2025) and SDSS (SDSS-Shen_2023) to identify and remove poorly performing for lags recovered by more complex methods.
3.2.2 PyROA
@@ 4 LITMUS Methodology
In this section, we introduce LITMUS, a new lag recovery package, based in python, making use of modern computational tools like jax (jax) and its ecosystem of flexible software like the statistical modelling framework numpyro (numpyro) and the GP modelling package tinyGP (cite_tinygp). LITMUS’s core feature is an algorithm for exploring and integrating the posterior that we call the Laplace Quadrature, which efficiently explores the parameter space to lags while also finding an estimate of the model evidence such that proper Bayesian null hypothesis testing is possible  [^2] Note that this method bears some similarity to a special case of the Integrated Nested Laplace Approximation (INLA Rue_2009_INLA; vanniekerk_2022_INLA). LITMUS also offers a flexible and modular statistical framework that allows it to be easily extended to new generative models, unlike existing tools which hard-code their statistical description of the AGN signals. We then present a new tool for RM that is more precise, mode correct and more robust than anything else in the literature while also being significantly faster and more broadly applicable.
@@ 4.1 The Laplace Quadrature
The core idea of the Laplace Quadrature is to side-step the rough geometry of the posterior along the lag-axis (e.g. see the bottom panel in Figure 2) by not trying to navigate over this geometry at all. Instead, a grid of test-lags are spaced along the lag-axis, and the shape of the posterior along all non-lag parameters is estimated with the Laplace approximation. In contrast to MCMC-like strategies which need to explore each axis of parameter space to marginalise over them, this Laplace approximation approach relies only on optimisation, for which the complexity scales much less severely with dimensionality of the parameter space. As such, the Laplace Quadrature approach can be applied up to markedly higher model dimensions without incurring an exorbitant computational burden.
@@
Gaussian distributions have easily evaluable integrals, and so the Laplace approximation also lets us easily calculate $Z_{i}$ for each slice:
@@
    Calculate the density $\mathcal{P}_{i}(\hat{\phi}^{i})$ and Hessian matrix $H_{i}(\phi=\hat{\phi})$ of the conditional distribution.
@@
A simplified example of this for a case where $\phi=\{\tau\}$ and $I=5$ is shown in Figure 4. Evenly spaced test lags are wasteful and error-prone, and so LITMUS uses a “grid smoothing” algorithm, outlined in Appendix A, to preferentially sample good-fit lags. For the optimisation to find $\hat{\phi}^{i}$ we use jaxopt’s implementation of the BFGS algorithm (Broyden_1970_BFGS01; Fletcher_1970_BFGS02; Goldfarb_1970_BFGS03; Shanno_1970_BFGS04), using also a pre-conditioning scheme described in Appendix B. Assuming that $\hat{\phi}$ is a smooth and continuous function of $\Delta t$, we use the solution $\hat{\phi}^{i}$ as the starting point for finding $\hat{\phi}^{i+1}$. To find the Hessian matrix, we use jax’s autodiff tool-set on the probability log-density functions provided for a particular model by numpyro.
@@ 4.2 Statistical Modelling
The Bayes factors comparing these models allows us to test the significance of a lag recovery. For an example of how these ratios can be interpreted, see Section 5.
@@ 4.2.1 Alternative Statistical Models
A useful feature of LITMUS is that, unlike existing tools which hard-code their statistical models, its modular design means that LITMUS can also fit alternate statistical descriptions of the signals, e.g. the null hypothesis models used to test the significance of lag recoveries. As an example, LITMUS also includes an alternative to the normal GP modelling but using an adjustable log-normal prior of the kind that would be recovered from an $R-L$ relationship, i.e. $\pi(\Delta t)\sim\mathcal{N}$.
This modularity also allows for more expressive models to account for noise or calibration uncertainty, for example:
  • Adding a “jitter” term to the response, $y_{r}(t)=y_{\mathrm{cont}}(t)\pm\sigma(t)$,
  • Error calibration scaling in the style of PyROA and CREAM, or
  • Modelling outlier observations with mixture model
are all easily implementable in LITMUS if there is a necessary use case.
LITMUS is easily extensible, any model that can expressed with numpyro can be fit by the Laplace and SVI quadrature methods. Including a smoothing kernel for the response, modelling the BLR holiday(Dehghanian_2019), and any number of other extensions are all possible with this same fitting tool.
4.2.2 The Constrained & Unconstrained Domain
Many Bayesian fitting algorithms encounter difficulties with hard boundaries on model priors, and perform best when parameters are “unconstrained”, i.e. with a domain of $\theta\in\mathbb{R},\;-\infty<\theta<\infty$. Such hard boundaries have infinite gradients in log-probability and other pathologies when paired with MCMC and SVI methods in RM, where convention is to use bounded uniform priors on all parameters, this affects all parameters. numpyro obviates these difficulties with a statistical slight of hand. Prior to fitting, all parameters with constrained priors are shifted with a change of variables into a new coordinate space in which they are unconstrained, e.g. for a Uniform distribution on $\theta\sim U_{a,b}(\theta)$ there’s a mapping $f:(a,b)\rightarrow\mathbb{R},\;\theta^{\prime}=f(\theta)$.
$P(\theta)=P(\theta)\times{J(\theta)},$ (19)
where $J(\theta)$ is the Jacobian of the transformation $\theta\rightarrow\theta^{\prime}$. The Laplace Quadrature and SVI Quadrature are all performed in this coordinate system by default, i.e. they assume that the posterior is approximately Gaussian in this unconstrained domain.  [^3] Strictly speaking all densities should be written in terms of $\phi^{\prime}$ and $\theta^{\prime}$ in the description of these algorithms, but we exclude this for simplicity of reading.
4.2.3 Convergence & Uncertainty Estimates
$\Delta\ln\left|Z_{i}\right|=-\nabla f^{T}H^{-1}\nabla f.$ (20)
Conveniently, this also has a natural interpretation in terms of the “closeness to the peak”. Noting that the Hessian is the negative inverse of the covariance matrix, this may also be written $-\Delta\phi^{T}H\Delta\phi=\Delta\phi^{T}C^{-1}\Delta\phi$, i.e. the square ‘number of standard deviations’ that the estimated solution is from the peak under the Gaussian approximation. This gives both a measure of the numerical uncertainty in $\ln\left|Z_{i}\right|$ and also a natural measure of “closeness” to the optimum.
$E_{\mathrm{int}}\approx\frac{1}{\sqrt{17}}\lvert Z-Z_{1/2}\rvert$ (21)
4.3 Alternate Algorithms
As well as the Laplace Quadrature, LITMUS includes other lag recovery algorithms for comparison. Firstly, it includes a jax based high speed implementation of ICCF. Secondly, it includes a variation of the Laplace Quadrature that uses SVI in place of the Laplace approximation, and finally it has an interface with the nested sampling (Skilling_2006_NS), specifically JAXNS (Albert_2020_JAXNS).
@@ 4.3.1 The SVI Quadrature
The Laplace Quadrature relies on Laplace approximation, which can be fragile against highly non-Gaussian posteriors. As an extension LITMUS also offers the SVI Quadrature, which has similar working but uses stochastic variational inference (SVI) to generate its Gaussian distribution approximation. While the Laplace approximation fits a Gaussian with a point-estimate based on the Hessian of the log-posterior at its maximum, SVI seeks to fit a Gaussian that is ‘most similar on average’ (see Figure 5 for a visual demonstration). This is codified by the Kullback–Leibler divergence (KL divergence), or “average log difference” between the Gaussian ‘surrogate distribution’ in Equation 15 and the true posterior density:
@@
The exact value of the $\mathrm{ELBO}$ is not easily evaluable, and is instead estimated from the summation in Equation 24. This is a necessarily stochastic measure as it relies on a finite set of random samples. As such maximising the $\mathrm{ELBO}$ makes use of stochastic optimisation. In LITMUS’s case, use numpyro’s ready-built SVI tool set and the stochastic optimiser ‘adam’ (kingma_2017_ADAM). Aside from using SVI to estimate the slice of the Gaussian slices and the $\mathrm{ELBO}$ to estimate the slice evidence, the SVI Quadrature operates identically to the Laplace Quadrature in Section 4.1.
Figure 5: A demonstration of the difference in the Laplace and SVI approximations, both attempting to emulate a Cauchy distribution (black solid line). The Laplace approximation (blue dotted line) creates a Gaussian that matches the curvature at the MAP of the true distribution, and in this case under-estimates the distribution everywhere else. The SVI approximation, here also fitting a Gaussian, instead tries to get as close as possible to the true distribution ‘on average’, and so over-estimates in the core region while balancing the impact on the evidence integral with under-estimates in the distribution’s tails.
@@ 5 Validation of Results
In this way we can use Bayes factors (ratios of the models’ Bayesian evidences) to test null hypotheses via model comparison. Here we formalise this process as asking two questions: is there a signal in the response, and if so does this signal demonstrate a significant lag? We consider a ”strong” result to be a Bayes factor of 100×100$\times$100 × or more, and a moderate result to be a factor of 10×10$\times$10 × or more. Using the Laplace Quadrature we find that we successfully identify the true lag where it exists, and our evidence calculations allows us to identify and exclude the spurious recoveries for the two false mocks at high SNR (left side of the bottom panel in figure 6 and offers reasonable caution where the answer is unclear at low SNR (bottom panel, right side).
Figure 6: A demonstration of LITMUS’s ability to use Bayes factors to distinguish if a signal is lag-carrying and/or contains structure. The top panels are mock light curves emulating the cadence and uncertainty of OzDES, both continuum (blue) and response (pink) signals. These mocks are broken into two cases: on the left, mocks with a low measurement uncertainty and easily observable slow fluctuations with a variational timescale signal of $\tau=200\mathrm{d}$, while the right panels show mocks with a higher measurement uncertainty and a more rapid variational timescale of $\tau=50\mathrm{d}$ that is difficult to distinguish from white noise with such a coarse observational cadence. The rows from top to bottom have the same continuum but different response signals. From top to bottom, the rows show a simulated reverberation response with a lag of $\Delta t=540\mathrm{d}$, a response that follows the same DRW but is decoupled such that there is no observable lag (equivalent to $\Delta t\rightarrow\infty$) and a response that is drawn purely from white noise (equivalent to a variational timescale of $\tau=0$). The bottom panel shows the differences in log-Bayes factors for each of the hypotheses in Table1, with orange bars being how strongly we can see structure in the response signals and blue bars being how well we can confirm the existence of a lag from this structure, with dotted lines indicating different significance levels in favour of accepting or rejecting these hypotheses. Red and yellow dots indicate the ground truth for each question: green means true, red means false. In the left examples we see that LITMUS has successfully recovered the truth for all tests at strong significance (all true positives and true negatives), and in the right examples we see that it even when the constraints are weak it gives no false positives.

### v2

@@ 3 Existing Methods
Reverberation mapping has progressed markedly over its decades long history, both in our understanding of the physics and of the best-practice in constraining the lags. Lag recovery methods can be broadly grouped into two categories: the Bayesian GP methods which make use of our understanding of the AGN signal statistics, and the non-GP methods which aim to be as agnostic as possible about the signal, using only flexible means of interpolation and measures of goodness of fit.
LITMUS’s main fitting methods belong to the GP-based class, and so in Section 3.1 we discuss this family and, in examine in particular its widely used exemplar JAVELIN and its limitations. We also discuss non-GP methods in Section 3.2, in particular the method used by PyCCF. In Section 3.3 we give a brief overview of methods that are not used by OzDES/ not directly compared against LITMUS in this paper.
3.1 Gaussian Process Methods / JAVELIN
The main focus of this paper is the class of GP methods that leverage our understanding of AGN signal statistics to work within the framework of a full Bayesian generative model. The core principle of these methods is to take the description of the AGN variability as a Gaussian process (discussed in detail in Section 2), for which there is a closed form and (somewhat) easily evaluable likelihood function.
@@ 3.1 Gaussian Process Methods / JAVELIN
JAVELIN is the longest standing of the GP fitting methods and has the distinction of being the most widely used. JAVELIN (JAVELIN-Zu_2010) is a successor to the FORTRAN-based SPEAR (Zu_2011_JAVELINpaper). It models the AGN light curves as a damped random walk (following the covariance functions outlined in Section 2) and models the response smoothing with a “top-hat” smoothing function:
@@
Noting that the exact form of smoothing tends to have little impact on the lag recovery in BLR RM (OZDES-Yu_2019) it is common-place when using JAVELIN to fix the smoothing scale, $b$, to some fixed value shorter than the observation cadence (e.g. OzDES-Penton_2021), a choice found to have little impact on lag recovery in OzDES-like observational surveys (OZDES-Yu_2019).
@@
Other GP-based lag recovery programs include CREAM (starkey_2017_creamerrorscaling) and MICA (Li_2016_MICA), discussed in more detail in Section 3.3. In this paper we focus our attention on JAVELIN’s AEIS fitting method, as JAVELIN is the primary OzDES lag recovery code and has a history as the most prevalent tool for BLR RM.
3.2 Non-Gaussian Process Methods / PyCCF
GP methods offer the most complete way to perform lag recovery, they are also more complicated and computationally arduous. Aside from costing sheer time to compute, this also opens room for numerical errors (see discussion of the aliasing mixing problem in Section 2.1). For this reason, the less rigorous but more exhaustive family of non-GP methods, which make as few assumptions about the underlying signal properties as possible, are still actively used in conjunction with or in preference over their more statistically precise GP cousins.
Such non-GP methods are valuable in that their lack of loose commitment towards a particular signal model lends them a flexibility and vagueness that absorbs our uncertainty about the AGN signal properties. Such models tend to also be numerically inexpensive, allowing exhaustive searches of their respective parameter spaces forcing their way past potential numerical challenges. Though LITMUS is primarily a GP method, we include these here for the sake of completeness and comparison.
@@ 3.2 Non-Gaussian Process Methods / PyCCF
The ICCF method has been found to agree with more rigorous models like JAVELIN to within statistical bounds, though with higher reported uncertainties (OZDES-Yu_2019). Owing to its numerical robustness and low computational overhead, it is still used as a validation tool by both OzDES (OzDES-Malik_2023; OzDES-Yu_2023; OzDES-Penton_2025) and SDSS (SDSS-Shen_2023) to identify and remove poorly performing for lags recovered by more complex methods. For comparison, LITMUS includes a jax accelerated implementation of the ICCF method in its array of fitting algorithms.
In their final RM data release, SDSS-Shen_2023 make use of the non-GP method PyROA (pyroa-ferus_2021), discussed in more detail in Section 3.3. In this paper we focus on a comparison with the ICCF method, which OzDES uses as a secondary method as a part of its validation and quality cuts.
3.3 Other Methods
Here we give a brief overview of three lag recovery methods that have been widely used outside of OzDES’s RM program: CREAM, MICA and PyROA.
CREAM
The GP-based fitting method CREAM (Starkey_2015_CREAM) is similar to JAVELIN in its modelling of the AGN light curve as a stochastic process and using Bayesian fitting, but differs in two key respects. Firstly, instead of fitting for the parameters of the observed signals it instead fits for physical properties of the AGN (e.g. mass, accretion rate etc.). Secondly it models the light curves by fitting the phase and amplitude of a series of Fourier components and, unlike JAVELIN, does not include the covariance of the observations in its likelihood function, instead using a $\chi^{2}$ goodness of fit loss function. A novel feature of CREAM is that it can be set to rescale the error bars of measurements from different telescopes as a part of the MCMC fitting in an internal calibration process (starkey_2017_creamerrorscaling). CREAM is used as the primary lag recovery method by some SDSS releases (SDSS-Grier_2017; CIV_Grier_2019), and is included in the reported results of their final data release (SDSS-Shen_2023).
MICA
Another GP-based fitting method from Li_2016_MICA is the code MICA. The main feature of this code is that it constructs the transfer function of Equation 5 from a sum of Gaussian profiles with positions and widths as free parameters that are marginalised over in the fitting process. In this way, instead of using a simplified approximate transfer function in the style of JAVELIN’s top-hat, MICA can allow for a more detailed accounting of the BLR’s geometry. MICA uses much of the same statistical approach as JAVELIN, but uses the Metropolis-Hastings algorithm (Metropolis_1953-MetHaste) in place of emcee.
PyROA
@@ 4 LITMUS Methodology
In this section, we introduce LITMUS , a new lag recovery package, based in python, making use of modern computational tools like jax (jax) and its ecosystem of flexible software like the statistical modelling framework numpyro (numpyro) and the GP modelling package tinyGP (cite_tinygp). LITMUS’s core feature is an algorithm for exploring and integrating the posterior that we call the Laplace Quadrature, which efficiently explores the parameter space to lags while also finding an estimate of the model evidence such that proper Bayesian null hypothesis testing is possible.  [^2] Note that this method bears some similarity to a special case of the Integrated Nested Laplace Approximation (INLA Rue_2009_INLA; vanniekerk_2022_INLA). LITMUS also offers a flexible and modular statistical framework that allows it to be easily extended to new generative models, unlike existing tools which hard-code their statistical description of the AGN signals. We then present a new tool for RM that is more precise, mode correct and more robust than anything else in the literature while also being significantly faster and more broadly applicable.
@@ 4.1 The Laplace Quadrature
The core idea of the Laplace Quadrature is to side-step the rough geometry of the posterior along the lag-axis (e.g. see the bottom panel in Figure 2) by not trying to navigate over this geometry at all. Instead, a grid of test-lags are spaced along the lag-axis, and the shape of the posterior along all non-lag parameters is estimated with the Laplace approximation (approximating a distribution as Gaussian by a second order expansion of its log-density). In contrast to MCMC-like strategies which need to explore each axis of parameter space to marginalise over them, this Laplace approximation approach relies only on optimisation, for which the complexity scales much less severely with dimensionality of the parameter space. As such, the Laplace Quadrature approach can be applied up to markedly higher model dimensions without incurring an exorbitant computational burden.
@@
As LITMUS is built in jax and numpyro, this Hessian matrix can be evaluated with the need for hand-calculated derivatives. Gaussian distributions have easily evaluable integrals, and so the Laplace approximation also lets us easily calculate $Z_{i}$ for each slice:
@@
[…]
@@
A simplified example of this for a case where $\phi=\{\tau\}$ and $I=5$ is shown in Figure 4. Evenly spaced test lags are wasteful and error-prone, and so LITMUS uses a “grid smoothing” algorithm, outlined in Appendix B, to preferentially sample good-fit lags. For the optimisation to find $\hat{\phi}^{i}$ we use jaxopt’s implementation of the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm (Broyden_1970_BFGS01; Fletcher_1970_BFGS02; Goldfarb_1970_BFGS03; Shanno_1970_BFGS04), using also a pre-conditioning scheme described in Appendix C. Assuming that $\hat{\phi}$ is a smooth and continuous function of $\Delta t$, we use the solution $\hat{\phi}^{i}$ as the starting point for finding $\hat{\phi}^{i+1}$. To find the Hessian matrix, we use jax’s autodiff tool-set on the probability log-density functions provided for a particular model by numpyro.
@@ 4.2 Statistical Modelling
The Bayes factors comparing these models allows us to test the significance of a lag recovery. For an example of how these ratios can be interpreted, see Section 4.4.
@@ 4.2.1 Alternative Statistical Models
A useful feature of LITMUS is that, unlike existing tools which hard-code their statistical models, its modular design means that LITMUS can also fit alternative statistical descriptions of the signals, e.g. the null hypothesis models used to test the significance of lag recoveries. As an example, LITMUS includes an alternative to the normal GP modelling but using an adjustable log-normal prior of the kind that would be recovered from an RLR-L relationship, i.e. $\pi(\Delta t)\sim\mathcal{N}$. Similar extensions can be made to incorporate the flexible transfer functions of MICA or the physically motivated parameterization and error calibration of CREAM, features that LITMUS’s initial release lacks.
This modularity also allows for more expressive models to probe physical phenomena, for example:
  • Using a summation of additional GP kernels to examine differences between the signal and the canonical DRW,
  • Modelling the BLR holiday (Dehghanian_2019), e.g. with a short-lived signal super-imposed on the GP signal, or
  • Testing models with time-varying lags to investigate changes in the BLR shape or scale
are all implementable in LITMUS if they can be expressed with a numpyro generative model, and can be fit by a suitable choice of its existing fitting methods.
4.3 Additional Algorithms
As well as the Laplace Quadrature, LITMUS includes other lag recovery algorithms for comparison. Firstly, it includes a jax based high speed implementation of ICCF. Secondly, it includes a variation of the Laplace Quadrature that uses stochastic variational inference (SVI) in place of the Laplace approximation, and finally it has an interface with the nested sampling (Skilling_2006_NS), specifically JAXNS (Albert_2020_JAXNS).
@@ 4.3.1 The SVI Quadrature
The Laplace Quadrature relies on Laplace approximation, which can be fragile against highly non-Gaussian posteriors. As an extension LITMUS also offers the SVI Quadrature, which has similar working but uses SVI to generate its Gaussian distribution approximation. While the Laplace approximation fits a Gaussian with a point-estimate based on the Hessian of the log-posterior at its maximum, SVI seeks to fit a Gaussian that is ‘most similar on average’ (see Figure 5 for a visual demonstration). This is codified by the Kullback–Leibler divergence (KL divergence), or “average log difference” between the Gaussian ‘surrogate distribution’ in Equation 14 and the true posterior density:
@@
The exact value of the $\mathrm{ELBO}$ is not easily evaluable, and is instead estimated from the summation in Equation 19. This is a necessarily stochastic measure as it relies on a finite set of random samples. As such maximising the $\mathrm{ELBO}$ makes use of stochastic optimisation. In LITMUS’s case, use numpyro’s ready-built SVI tool set and the stochastic optimiser ‘adam’ (kingma_2017_ADAM). Aside from using SVI to estimate the Gaussian slices and the $\mathrm{ELBO}$ to estimate the slice evidence, the SVI Quadrature operates identically to the Laplace Quadrature in Section 4.1.
Figure 5: A demonstration of the difference in the Laplace and SVI approximations, both attempting to emulate a Cauchy distribution (black solid line). The Laplace approximation (blue dotted line) creates a Gaussian that matches the curvature at the MAP of the true distribution, and in this case under-estimates the distribution everywhere else. The SVI approximation, here also fitting a Gaussian, instead tries to get as close as possible to the true distribution ‘on average’, and so under-estimates in the core region while balancing the impact on the evidence integral with over-estimates in the distribution’s tails.
@@ 4.4 Hypothesis Testing
In this way we can use Bayes factors (ratios of the models’ Bayesian evidences) to test null hypotheses via model comparison. Here we formalise this process as asking two questions: is there a signal in the response, and if so does this signal demonstrate a significant lag? We consider a ”strong” result to be a Bayes factor of $100\times$ or more, and a moderate result to be a factor of $10\times$ or more. Using the Laplace Quadrature we find that we successfully identify the true lag where it exists, and our evidence calculations allow us to identify and exclude the spurious recoveries for the two false mocks at high SNR.
Figure 6 contains a demonstration of LITMUS’s ability to use Bayes factors to distinguish if a signal is lag-carrying and/or contains structure. We simulate mock light curves emulating the cadence and uncertainty of OzDES in two cases: mocks with a low measurement uncertainty and easily observable slow fluctuations with a variational timescale signal of $\tau=200\mathrm{d}$, and mocks with a higher measurement uncertainty and a more rapid variational timescale of $\tau=50\mathrm{d}$ that is difficult to distinguish from white noise with such a coarse observational cadence. For each case we simulate a single continuum but three different response signals: a simulated reverberation response with a lag of $\Delta t=540\mathrm{d}$, a response that follows the same DRW but is decoupled such that there is no observable lag (equivalent to $\Delta t\rightarrow\infty$) and a response that is drawn purely from white noise (equivalent to a variational timescale of $\tau=0$). Using the Laplace Quadrature’s evidence integrals, LITMUS can successfully recover the truth for all tests at strong significance (all true positives and true negatives) in the high SNR case, and returns no false positives even when the constraints are weak.
Figure 6: Mocks and evidence ratios for the demonstrative mock signals in the body of the text. The top panels shows the mock light curves for continuum (blue) and response (pink) signals. The left and right columns correspond to high and low SNR, while the rows from top to bottom show the mocks for a coupled response at a lag of $\Delta t=540\mathrm{d}$, a decoupled response and a pure white noise response. The bottom panel shows the differences in log-Bayes factors for each of the hypotheses in Table1, with orange bars being how strongly we can see structure in the response signals and blue bars being how well we can confirm the existence of a lag from this structure, with dotted lines indicating different significance levels in favour of accepting or rejecting these hypotheses. Red and green markers dots indicate the ground truth for each question: green circles mean true, red squares mean false.